OpenStudy (anonymous):

Convert -3R - 4U to polar form

OpenStudy (anonymous):

what you mean polar form? complex number or polar cordinates?

OpenStudy (anonymous):

polar coordinates

OpenStudy (anonymous):

is that yu are given ?

OpenStudy (anonymous):

all*

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[x = r \cos \theta \leftarrow \rightarrow y =r \sin \theta\]\[\theta = \tan^{-1} (y/x)\] Does this help ?

OpenStudy (anonymous):

not really :( sorry!

OpenStudy (anonymous):

hmm... ok. so are you assuming R is y and U is x in this case then ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok so rearranging we have: R = x = r cos (@) U = y = r sin (@) hence, -3R - 4U -3 (r cos(@)) - 4 (r sin(@)) where @ = tan^(-1) (u/r) = tan^-1 [(r sin@)/(r cos@)] r = sqrt { r^2 + u^2 } can you solve it now?

OpenStudy (anonymous):

yes it think so thanks alot!!!!

OpenStudy (anonymous):

np